Ticket Number: HEA-NOC/20040715-2 Ticket Status: UPDATE
Ticket Type: scheduled Resolver: HEA-NOC
Ticket Opened: 20040715 09:30 UTC+1 Problem Start: 20040715 12:00 UTC+1
Ticket Update: 20040715 12:37 UTC+1 Problem End: 20040715 12:30 UTC+1
Site/line: MX backup, mail
There will be a brief outage (roughly 10 mins) on mail.heanet.ie to install
new disks and memory.
20040715 09:22 CMC Opening ticket
20040715 12:36 CMC mail.heanet.ie upgraded, however disks are taking
slightly longer than expected to syncronise. Outage expected to last a
further 10 minutes.
This ticket can be monitored at http://www.hea.net/tickets/20040715-2